Chapter 1.4

The Concept of Independence. The Bernoulli Scheme

[7/29%]
Show
LaTeX
Problem 1.4.1

Balls are drawn from a box containing black and white balls. Let the event AkA_{k} mean that a white ball is drawn on the kk-th step. Prove that the events A1,…,AnA_{1}, \ldots , A_{n}

  1. are mutually independent, if the balls are drawn with replacement;

  2. are dependent, if the balls are drawn without replacement.

?
Problem 1.4.2

Two players carry out a series of independent trials. In each trial, player AA throws 3 dice, while player BB throws 2 dice simultaneously with player AA and independently of him. They carry out these trials in succession until the first occurrence of a "six" on at least one of the dice. Find the probabilities of the following events:

  1. C={ the "six" first appeared for player B, and not for A}\mathcal{C}=\left\{ \text{ the "six" first appeared for player }B\text{, and not for }A\right\}

  2. D={ the "six" first appeared simultaneously for A and B}\mathcal{D}=\left\{ \text{ the "six" first appeared simultaneously for }A\text{ and }B\right\}.

?
Problem 1.4.3

Let A,B,CA, B, C be pairwise independent equiprobable events such that A∩B∩C=∅A \cap B \cap C=\varnothing. Find the maximum possible value of P(A)\mathbb {P}\left(A\right).

?
Problem 1.4.4

A set SS of nn elements is given. Three subsets A,B,CA, B, C are chosen from it randomly and independently. Each random subset is formed as follows: each element of the set SS, independently of the others, is included in the subset with probability pp, and is not included with probability (1−p)(1-p) (this process is repeated three times: first for the set AA, then for BB, and finally for C)C). Find the probability of the event D={A∩B⊆C⊆A∪B}D=\left\{ A \cap B \subseteq C \subseteq A \cup B\right\}.

?
Problem 1.4.5

A social network user views the wall of one of his 7 friends each day with probability pi,i=1,…,7p_{i}, i=1, \ldots , 7; ν\nu is the number of the first week during which he views the walls of all seven of his friends. Find the probability that on the Monday of week ν\nu he views the wall of his first friend.

?
Problem 1.4.6

The edges of the complete graph KnK_{n} on nn vertices are colored independently of each other, with equal probability 1k\frac{1}{k}, in one of kk colors. Let VV be the set of vertices of the graph KnK_{n}, and S⊂VS \subset V. Denote by ASA_{S} the following event: AS={ all edges of Kn both of whose vertices belong to S are colored the same color }A_{S}=\left\{ \text{ all edges of }K_{n}\text{ both of whose vertices belong to }S\text{ are colored the same color }\right\}. Under what conditions on the mutual arrangement of the subsets S,T⊂VS, T \subset V are the events ASA_{S} and ATA_{T} independent?

?
Problem 1.4.7

Let A1,…,AnA_{1}, \ldots , A_{n} be independent events. Prove that

P(⋃i=1nAi)=1−∏i=1nP(Ai‾) \mathbb {P}\left(\bigcup _{i=1}^{n} A_{i}\right)=1-\prod _{i=1}^{n} \mathbb {P}\left(\overline{A_{i}}\right)
?